Which is the equation of a circle that passes through(2,2) and is centered at(5,6)?
Answer by lwsshak3(11628)

Which is the equation of a circle that passes through(2,2) and is centered at(5,6)?
**
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
Solving for radius using coordinates of given point (2,2)
(2-5)^2+(2-6)^2=r^2
9+16=r^2
r^2=25
Equation of circle:
(x-5)^2+(y-6)^2=25